OpenStudy (anonymous):

The circumference of a circular painting is 21.98 feet. What is the diameter of the painting?

OpenStudy (anonymous):

6.996, or 7 ft (I think) :P

OpenStudy (jdoe0001):

\(\bf \textit{Circumference of a Circle}=2\pi r\qquad r=radius\implies \cfrac{diameter}{2}\\ \quad \\ 21.98=2\pi r\implies \cfrac{21.98}{2\pi}=r\qquad diameter = 2\times r\)